### v3

@@
Rotation matters for the life of a star. It causes a multitude of dynamical phenomena in the stellar interior during a star’s evolution and its effects accumulate until the star dies. All stars rotate at some level but most of those born with a mass above about 1.3 times the mass of the Sun rotate rapidly during more than 90% of their nuclear lifetime. Internal rotation guides the angular momentum and chemical element transport throughout the stellar interior. These transport processes change over time as the star evolves. The cumulative effects of stellar rotation and its induced transport processes determine the helium content of the core by the time it exhausts its hydrogen isotopes. The amount of helium at that stage also guides the heavy element yields at the end of the star’s life.
A proper theory of stellar evolution and any realistic models for the chemical enrichment of galaxies, must be based on observational calibrations of stellar rotation and of the induced transport processes. In the last few years, asteroseismology offers such calibrations, for single and binary stars. We review the current status of asteroseismic modelling of rotating stars for different stellar mass regimes, in an accessible way for the non-expert. While doing so, we describe exciting opportunities sparked by asteroseismology for various domains in astrophysics, touching upon topics from exoplanetary science to galactic structure and evolution and on to gravitational wave physics. Along the way, we provide ample sneak-previews for future ‘industrialised’ applications of asteroseismology to slow and rapid rotators, from exploitation of combined Kepler, TESS, PLATO, Gaia, and spectroscopy surveys.
We end the review with a list of take away messages and achievements of asteroseismology, which are of relevance for many fields of astrophysics.
Asteroseismic Modelling
of Fast Rotators and its
Opportunities for Astrophysics
@@
Asteroseismology – Waves – Binary Stars – Stars: Rotation – Stars: Interiors – Stars: evolution – Stars: oscillations (including pulsations)– Stars: magnetic field – Gravitational Waves – Convection – Hydrodynamics – Methods: data analysis – Methods: statistical – Surveys – Sun: interior – Sun: helioseismology – Binaries: eclipsing – Stars: subdwarfs – Stars: supergiants – Galaxy: clusters: general – Stars: black holes – Stars: neutron – Binaries (including multiple): close – Blue stragglers – Stars: emission-line, Be1 Asteroseismology: a fountain of opportunities for astrophysics
@@
Asteroseismology is the science of probing stellar interiors by modelling detected oscillation modes of stars (see Aerts et al. 2010, for a comprehensive monograph). It has been an established research field within stellar astrophysics for more than a decade now. Rapid progress in numerous successful applications of this technique has occurred thanks to major advances in the observational aspects brought by space telescopes. This is mainly thanks to space missions dedicated to the collection of uninterrupted high-cadence high-precision photometric light curves, such as the past MOST (Walker et al. 2003), CoRoT (Auvergne et al. 2009), and Kepler (Koch et al. 2010) satellites and the currently operational TESS mission (Ricker et al. 2015). The oscillation frequencies deduced from such time-series data are independent of stellar models as they are derived directly from the Fourier transforms of the light curves. The mode frequencies are typically one to several orders of magnitude more precise than classical observables used to evaluate stellar structure and evolution models in Hertzsprung-Russell (HRD) or colour-magnitude (CMD) diagrams (cf. Table 1 in Aerts et al. 2019). Moreover, oscillation mode frequencies are influenced by the physical and chemical conditions of the matter in the part of the stellar interior where the modes have their dominant probing power. It is then no surprise that the exploitation of seismic observables determined directly by the deep interior of stars brings a revolution for astrophysics (Aerts 2021). Now that the field of asteroseismology is mature and well established, we enter its ‘industrial revolution’ thanks to the large surveys brought by TESS and the PLATO mission currently under construction (Rauer et al. 2014, see Rauer et al. (submitted) for the latest updates).
@@
Figure 1 represents a fountain of opportunities fed by asteroseismology, as a non-exhaustive preview of the themes touched upon in this review. Indeed, asteroseismology currently plays an important role in the progress of each of these topics of modern astrophysics. This fountain highlights the major capacities of asteroseismology by uncovering the detailed properties of stars, all the way from their central core to their surface. Thanks to this unique capability, asteroseismology brings novel and crucial ways to improve stellar structure and evolution theory, impacting all research in astrophysics relying on stellar models. For each of the topics shown in Fig. 1, we highlight the role of asteroseismic input, with emphasis on improvements achieved or to be expected in the not too distant future.
@@
The extensive general and observational papers by Aerts (2021) and Kurtz (2022), respectively, already offer comprehensive reviews on the history, beginnings, methodology, and numerous applications of asteroseismology in the current space era, written for a broad readership. Here we offer new insights and more recent applications for the general A&A reader, which are not yet covered or too recent to have been in these two comprehensive reviews. In particular, we focus this review on applications of asteroseismology to rapidly rotating stars and discuss why and how they differ from asteroseismology of the Sun-like stars. Such stars are slow rotators during the largest part of their life because they are subject to magnetic braking during their initial life phases.
@@
The character of an oscillation mode is mainly determined by its dominant restoring force(s). Examples are the pressure gradient for pressure (p) modes, also known as acoustic modes, the buoyancy force of Archimedes for gravity (g) modes, or the Coriolis force for inertial modes. In the absence of rotation, the family of spheroidal modes is categorised into three branches: the high-frequency p modes and the low-frequency g modes are separated by the so-called fundamental (f) mode, which corresponds to surface gravity waves. This simple picture becomes more complex in the presence of rotation. We come back to the different types of modes, which probe different layers of the star, in Sect. 3. In this review, we will be highlighting modelling results based on identified oscillation modes probing the deep internal physics of observed rotating stars. A key motivation for this emphasis is the major impact offered by the interpretation of so-called gravito-inertial oscillation modes. Such modes allow us to improve the theory of stellar evolution in the presence of rapid rotation. Gravito-inertial asteroseismology only saw its birth a few years ago and has not been covered extensively in previous review papers, while it is of major importance for stellar evolution and chemical yields.
@@
Rotation has a major impact during the longest nuclear life phase of all the single and binary stars of intermediate and high mass (Maeder 2009). In that sense, the internal rotation during that phase is a main driver for the chemical evolution of galaxies (Karakas & Lattanzio 2014; Kobayashi et al. 2020). It has to be kept in mind that the effects of rotation, notably the transport and mixing phenomena it induces, are cumulative throughout the evolution of the star. Rotationally induced internal mixing and angular momentum transport in stars born with a convective core and a radiative envelope leave strong fingerprints during the initial 90% of their life, that is, while they are burning hydrogen into helium in their rotating core.
@@
Throughout this so-called main-sequence phase the cumulative effect of rotationally induced or affected transport processes may lead to the production of far more helium compared to the case where rotation is not active or only mildly so. The efficiency of the heavy element production beyond the main sequence is largely determined by the amount of helium available as fuel in the final 10% of stellar life (Kippenhahn et al. 2013; Salaris & Cassisi 2017). During that short and complicated end phase of stellar evolution, stars are slowed down tremendously as they expand and lose angular momentum due to winds, outflows, and explosions (for those born with a mass above about eight solar masses). In this way, stars enrich their surrounding interstellar medium near the end of their life. To assess the amount and kind of processed nuclear material expelled by dying stars, it is essential to measure the internal rotation, mixing, and the masses of the growing helium and carbon-oxygen cores during the stages of central hydrogen and helium burning. Nowadays, asteroseismology is able to measure these helium and carbon-oxygen cores, as relics of the accumulated effect of the internal rotation and mixing throughout the two longest nuclear life phases of stars.
@@
Non-radial oscillations in fast rotators provide critical and new calibrations of the internal rotation and magnetic fields to stellar evolution theory. Aside from this basic role of asteroseismology, it also offers guides to improve 3-dimensional (3D) magneto-hydrodynamical (MHD) simulations of core-collapse supernovae (e.g. Hammer et al. 2010; Wongwathanarat et al. 2015; Lentz et al. 2015; Summa et al. 2016; Müller et al. 2017; Ott et al. 2018; O’Connor & Couch 2018; Andresen et al. 2019; Burrows et al. 2019; Varma et al. 2023) and to calibrate binary population synthesis studies. These guide predictions for gravitational wave emission from merging neutron star and black hole binaries (e.g., de Mink & Belczynski 2015; de Mink & Mandel 2016; Marchant et al. 2016; Belczynski et al. 2016; Farr et al. 2017; Belczynski et al. 2020; Laplace et al. 2020; Marchant et al. 2020, 2021; Landry & Read 2021; Schneider et al. 2021, 2023; Mezzacappa et al. 2023; Wong et al. 2023; Agrawal et al. 2023; Jiang et al. 2023; Cheng et al. 2023). No such simulation studies available today have been done with asteroseismically calibrated internal rotation and magnetic field profiles of massive stars. Major future endeavours are therefore to be anticipated on the bridge between asteroseismology of massive close binaries and multi-messenger astronomy. Asteroseismology is an essential approach to bring a solid observational foundation to the interpretation of gravitational wave detections of merging compact binaries (cf. Fig. 1).
@@ 2 Existing asteroseismology reviews and the aspect of binarity
Aside from the recent reviews by Aerts (2021) and Kurtz (2022) mentioned already, several earlier asteroseismology review papers exist. These were published during the past decade and are dedicated to particular types of pulsators, following the tremendous and rapid progress delivered by the 4-yr light curves assembled by the nominal NASA Kepler space mission (Koch et al. 2010) and its refurbished version K2 (Howell et al. 2014). These previous manuscripts offer comprehensive overviews of the impact of asteroseismology on astrophysics for specific types of stars, notably slow rotators. We shall not repeat their content here, but do take the opportunity to present some of the latest asteroseismology results for slow rotators not yet covered in these available reviews. In doing so we touch upon a few topics that stand out because of their unique potential for impact in future astrophysics research, notably internal magnetism of stars.
So far, the playground of asteroseismology, that is, the range in the birth mass of stars with identified non-radial oscillation modes, covers from about 0.75 M for the K5V star Indi A (Lundkvist et al. 2024) to as high as about 25 M for the O9V Cep star HD 46202 (Briquet et al. 2011). However, by far most efforts in asteroseismic modelling have been concentrated on low-mass stars, from their birth as dwarfs through their evolved stages as subgiants and red giants, or as subdwarfs due to loss of their hydrogen envelope, and all the way to their end-of-life as cooling compact remnant white dwarfs (see the reviews by Chaplin & Miglio 2013; Hekker & Christensen-Dalsgaard 2017; Córsico et al. 2019; García & Ballot 2019; Giammichele et al. 2022). Multiple reasons for this focus on stars born with a low mass (defined here as stars with a mass in the range 0.7 M M) are relevant. Indeed, low-mass stars
@@
constitute the dominant population in our Milky Way galaxy, allowing for the study of its history, archaeology, and current structure. This is nowadays being done from the multitude of Gaia data (Gaia Collaboration et al. 2016, 2023b) and ongoing or near-future dedicated all-sky survey high-resolution spectroscopy (Kollmeier et al. 2017; Pinsonneault et al. 2018; Bensby et al. 2019; Jin et al. 2023);
@@
reveal high-frequency solar-like oscillations excited stochastically by the turbulent convection in their envelopes throughout almost their entire nuclear burning life. Such oscillations obey so-called scaling relations already summarised by Kjeldsen & Bedding (1995) long before the era of space asteroseismology. These scaling relations make it easy to deduce the mass, radius, and age from radial, dipole, and quadrupole modes, with high precisions of a few percent. This is achieved by comparing the observed properties of detected low-degree global oscillation modes with those of the Sun (e.g. García & Ballot 2019, for an extensive review). It has the major advantage that one can use the heritage from helioseismology (Christensen-Dalsgaard 2002, 2021) and transfer it to asteroseismic applications to solar-like oscillations;
@@
cover the most important parameter space of exoplanet host stars, including those with multiple Earth-like rocky planets in the habitable zone;
@@
have a relatively modest binary and multiplicity fraction compared to stars of higher mass (cf. Fig. 2);
@@
are all slow rotators from early in their life as they are subject to efficient magnetic braking during their core hydrogen burning phase. As we will detail in the next sections, this implies that the impact of rotation on the modelling of their oscillations is in the easiest regime of the frequency domain. We come back to this important aspect throughout this review.
@@
None of these five aspects is valid for moderate to fast rotators of intermediate to high mass. We define stars born with 1.3 M M as intermediate-mass stars and those with M as high-mass stars. For the pulsators in these two mass regimes we do not have a main-sequence calibrator like the Sun, whose helioseismology (Christensen-Dalsgaard 2002) has been essential to guide similar applications to low-mass stars.
@@
Though a minority compared to their low-mass colleagues, stars born with M deliver a major fraction of the energy and heavy elements to galaxies and to the Universe as a whole. For this reason, and given the fast progress of asteroseismology for this under-represented category of fast rotators, we dedicate most of this review to asteroseismic applications of stars born with intermediate or high mass. In the next section we clarify what we mean by moderate and fast rotation.
@@
Fig. 2 shows the binary fraction for stars and brown dwarfs. This figure makes it clear that we cannot ignore stellar multiplicity in a review focusing on asteroseismology of intermediate- and high-mass stars. In particular, close binarity dominates the evolution of the most massive stars (Sana et al. 2012, 2013, 2014). While this has been firmly established observationally for a decade, it is recognised that binary evolution theory still suffers significantly from the major uncertainties occurring in the theory of single star evolution (Marchant & Bodensteiner 2024). Hence this must be given the highest priority and asteroseismology is the optimal way to do so (Aerts 2021). We will touch upon the various aspects of binarity and multiplicity for asteroseismic studies in Sect. 5 and Sect. 6. The class of subdwarfs is discussed briefly in Sect. 7. It deserves special attention in this respect, as the binary fraction is high (Han et al. 2003; Hu et al. 2008; Vos et al. 2017, 2019). In fact, all subdwarfs originate from binary interactions because the seemingly single class members can only be understood in terms of a common envelope phase resulting in a merger (Heber 2009). Due to the efficient loss of their envelope, the core-helium burning subdwarfs are of low mass in their current phase of evolution (about 0.5 M). Those with oscillations reveal high-frequency modes with periods much shorter than their rotation and/or orbital period (see Lynas-Gray 2021, for a review). We also briefly touch upon merger seismology for higher mass objects, offering a way to hunt for such stars asteroseismically. Section 8 is dedicated to future outlooks, putting emphasis on Gaia’s role in asteroseismology.
@@
Rather than trying to be exhaustive in reviewing the progress of asteroseismology, this paper takes the approach of offering the non-expert reader highlights from the recent literature based on prominent case studies. This way, we illustrate what can be achieved with current and future data and theoretical developments. The plan of the paper is as follows. We provide an accessible introduction to asteroseismology in Sect. 3, touching upon the wave equations representing stellar oscillations and the accompanying frequency regimes. Section 4 covers recent applications of asteroseismology of single stars, starting with the Sun and gradually increasing the level of complexity in the mode computations as the ratio of the oscillation mode periods to the rotation period increases. In Sect. 5 we highlight how model-independent dynamical masses and radii help asteroseismology, before moving on to tidal asteroseismology in Sect. 6 and subdwarf and merger seismology in Sect. 7. We end the paper with an extensive outlook for the bright future of asteroseismic applications to stars of intermediate and high mass from combined Gaia data, survey spectroscopy, and ongoing TESS or future PLATO space photometry. A pertinent need for new theory and better modelling tools occurs, in order to interpret the oscillations of fast rotators and close binaries up to current measurement precisions.

### v4

@@
Asteroseismic modelling of fast rotators and its opportunities for astrophysics
Rotation matters for the life of a star. It causes a multitude of dynamical phenomena in the stellar interior during a star’s evolution, and its effects accumulate until the star dies. All stars rotate at some level, but most of those born with a mass higher than 1.3 times the mass of the Sun rotate rapidly during more than 90% of their nuclear lifetime. Internal rotation guides the angular momentum and chemical element transport throughout the stellar interior. These transport processes change over time as the star evolves. The cumulative effects of stellar rotation and its induced transport processes determine the helium content of the core by the time it exhausts its hydrogen isotopes. The amount of helium at that stage also guides the heavy element yields by the end of the star’s life.
A proper theory of stellar evolution and any realistic models for the chemical enrichment of galaxies must be based on observational calibrations of stellar rotation and of the induced transport processes. In the last few years, asteroseismology offers such calibrations for single and binary stars. We review the current status of asteroseismic modelling of rotating stars for different stellar mass regimes in an accessible way for the non-expert. While doing so, we describe exciting opportunities sparked by asteroseismology for various domains in astrophysics, touching upon topics such as exoplanetary science, galactic structure and evolution, and gravitational wave physics to mention just a few. Along the way we provide ample sneak-previews for future ‘industrialised’ applications of asteroseismology to slow and rapid rotators from the exploitation of combined Kepler, Transiting Exoplanet Survey Satellite (TESS), PLAnetary Transits and Oscillations of stars (PLATO), Gaia, and ground-based spectroscopic and multi-colour photometric surveys.
We end the review with a list of takeaway messages and achievements of asteroseismology that are of relevance for many fields of astrophysics.
@@
Asteroseismology – Waves – Binary Stars – Stars: Rotation – Stars: Interiors – Stars: evolution – Stars: oscillations (including pulsations)– Stars: magnetic field – Gravitational Waves – Convection – Hydrodynamics – Methods: data analysis – Methods: statistical – Surveys – Sun: interior – Sun: helioseismology – Binaries: eclipsing – Stars: subdwarfs – Stars: supergiants – (Galaxy:) open clusters and associations: general – Stars: black holes – Stars: neutron – Binaries (including multiple): close – Blue stragglers – Stars: emission-line, Be1 Asteroseismology: A fountain of opportunities for astrophysics
@@
Asteroseismology is the science of probing stellar interiors by modelling detected oscillation modes of stars (see Aerts et al. 2010, for a comprehensive monograph). It has been an established research field within stellar astrophysics for more than a decade. Rapid progress in numerous successful applications of this technique has occurred due to major advances in the observational aspects brought by space telescopes. This is mainly thanks to space missions dedicated to the collection of uninterrupted high-cadence high-precision photometric light curves, such as the past satellites Microvariability and Oscillations of STars (MOST, Walker et al. 2003), Convection, Rotation, and planetary Transits (CoRoT, Auvergne et al. 2009), and Kepler (Koch et al. 2010), and the currently operational Transiting Exoplanet Survey Satellite telescope (TESS, Ricker et al. 2015). The oscillation frequencies deduced from such time series data are independent of stellar models as they are derived directly from the Fourier transforms of the light curves. The mode frequencies are typically one to several orders of magnitude more precise than the classical observables used to evaluate stellar structure and evolution models in Hertzsprung-Russell diagram (HRD) or colour-magnitude diagram (CMD) (see Table 1 in Aerts et al. 2019). Moreover, oscillation mode frequencies are influenced by the physical and chemical conditions of the matter in the part of the stellar interior where the modes have their dominant probing power. It is then no surprise that the exploitation of seismic observables determined directly by the deep interior of stars brings a revolution for astrophysics (Aerts 2021). Now that the field of asteroseismology is mature and well established, we enter its ‘industrial revolution’ thanks to the large surveys brought by TESS and the PLAnetary Transits and Oscillations of stars mission currently under construction (PLATO, Rauer et al. 2024).
@@
Figure 1 represents a fountain of opportunities fed by asteroseismology. It provides a non-exhaustive preview of the themes touched upon in this review. Asteroseismology currently plays an important role in the progress of each of these topics of modern astrophysics. This fountain highlights the major capabilities of asteroseismology by uncovering the detailed properties of stars, all the way from their central core to their surface. Thanks to this unique capability, asteroseismology brings novel and crucial ways to improve stellar structure and evolution theory, impacting all research in astrophysics that relies on stellar models. For each of the topics shown in Fig. 1, we highlight the role of asteroseismic input, with emphasis on the improvements achieved or to be expected in the not-too-distant future.
@@
The extensive general and observational papers by Aerts (2021) and Kurtz (2022), respectively, already offer comprehensive reviews on the history, beginnings, methodology, and numerous applications of asteroseismology in the current space era, written for a broad readership. Here we offer new insights and more recent applications for the general A&A reader that have not yet been covered or are too recent to have been included in these two comprehensive reviews. In particular, we focus this review on applications of asteroseismology to rapidly rotating stars, and discuss why and how they differ from the asteroseismology of the Sun-like stars. Such stars are slow rotators during most of their life because they are subject to magnetic braking during their initial life phases.
@@
The character of an oscillation mode is mainly determined by its dominant restoring force(s). Examples are the pressure gradient for pressure (p) modes, also known as acoustic modes; the buoyancy force of Archimedes for gravity (g) modes; and the Coriolis force for inertial modes. In the absence of rotation, the family of spheroidal modes is categorised into three branches: the high-frequency p modes and the low-frequency g modes are separated by the fundamental (f) mode, which corresponds to surface gravity waves. This simple picture becomes more complex in the presence of rotation. We come back to the different types of modes, which probe different layers of the star, in Sect. 3. In this review we highlight the modelling results based on identified oscillation modes probing the deep internal physics of observed rotating stars. A key motivation for this emphasis is the major impact offered by the interpretation of gravito-inertial oscillation modes. These modes allow us to improve the theory of stellar evolution in the presence of rapid rotation. Gravito-inertial asteroseismology only came into being a few years ago, and has not been covered extensively in previous review papers, while it is of major importance for calibrating stellar evolution and chemical yield computations (e.g. Hirschi et al. 2004, 2005; Kaiser et al. 2020; Brinkman et al. 2024).
@@
Rotation has a major impact during the longest nuclear life phase of all the single and binary stars of intermediate and high mass (Maeder 2009). In that sense, the internal rotation during that phase is an important driver for the chemical evolution of galaxies (Karakas & Lattanzio 2014; Kobayashi et al. 2020). It should be kept in mind that the effects of rotation, notably the transport and mixing phenomena it induces, are cumulative throughout the evolution of the star. Rotationally induced internal mixing and angular momentum transport in stars born with a convective core and a radiative envelope leave strong fingerprints during the initial 90% of their life, that is, while they are burning hydrogen into helium in their rotating core.
@@
Throughout this main-sequence phase, the cumulative effect of rotationally induced or affected transport processes may lead to the production of far more helium compared to the case where rotation is not active or only mildly so. The efficiency of the heavy element production beyond the main sequence is largely determined by the amount of helium available as fuel in the final 10% of stellar life (Kippenhahn et al. 2013; Salaris & Cassisi 2017). During that short and complicated end phase of stellar evolution, stars are slowed down tremendously as they expand and lose angular momentum due to winds, outflows, and explosions (for those born with a mass above about eight solar masses). In this way, stars enrich their surrounding interstellar medium near the end of their life. To assess the amount and kind of processed nuclear material expelled by dying stars, it is essential to measure the internal rotation, mixing, and the masses of the growing helium and carbon-oxygen cores during the stages of central hydrogen and helium burning. Asteroseismology is able to measure these helium and carbon-oxygen cores as relics of the accumulated effect of the internal rotation and mixing throughout the two longest nuclear life phases of stars.
@@
Non-radial oscillations in fast rotators provide critical and new calibrations of the internal rotation and magnetic fields to stellar evolution theory. Aside from this basic role of asteroseismology, it also offers guides to improve three-dimensional (3D) magnetohydrodynamical (MHD) simulations of core-collapse supernovae (e.g. Hammer et al. 2010; Wongwathanarat et al. 2015; Lentz et al. 2015; Summa et al. 2016; Müller et al. 2017; Ott et al. 2018; O’Connor & Couch 2018; Andresen et al. 2019; Burrows et al. 2019; Varma et al. 2023) and to calibrate binary population synthesis studies. These two topics, among others, guide predictions for gravitational wave emission from merging neutron star and black hole binaries (e.g. de Mink & Belczynski 2015; de Mink & Mandel 2016; Marchant et al. 2016; Belczynski et al. 2016; Farr et al. 2017; Belczynski et al. 2020; Laplace et al. 2020; Marchant et al. 2020, 2021; Landry & Read 2021; Schneider et al. 2021, 2023; Mezzacappa et al. 2023; Wong et al. 2023; Agrawal et al. 2023; Jiang et al. 2023; Cheng et al. 2023). No such studies available today have been done with asteroseismically calibrated internal rotation and magnetic field profiles of massive stars. Major future endeavours are therefore to be anticipated to study the bridge between asteroseismology of massive close binaries and multi-messenger astronomy. Asteroseismology is an essential approach to bring a solid observational foundation to the interpretation of gravitational wave detections in terms of the progenitors of their merging compact binaries (see Fig. 1).
@@ 2 Existing asteroseismology reviews and the aspect of binarity
Aside from the recent reviews by Aerts (2021) and Kurtz (2022) mentioned already, several earlier asteroseismology review papers exist. These works were published during the past decade and are dedicated to particular types of pulsators, following the tremendous and rapid progress delivered by the 4 yr light curves assembled by the nominal NASA Kepler space mission (Koch et al. 2010) and its refurbished version K2 (Howell et al. 2014). These previous manuscripts offer comprehensive overviews of the impact of asteroseismology on astrophysics for specific types of stars, notably slow rotators. We do not repeat their content here, but do take the opportunity to present some of the latest asteroseismology results for slow rotators not yet covered in the available reviews. In doing so we touch upon a few topics that stand out because of their unique potential for impact in future astrophysics research, notably the internal magnetism of stars.
To date, the playground of asteroseismology (i.e. the range in birth masses of stars with identified non-radial oscillation modes) is from about 0.75 M⊙ for the K5V star Indi A (Lundkvist et al. 2024) to as high as about 25 M⊙ for the O9V Cep star HD 46202 (Briquet et al. 2011). However, by far the most efforts in asteroseismic modelling have been concentrated on low-mass stars, from their birth as dwarfs through their evolved stages as sub-giants and red giants, or as sub-dwarfs due to loss of their hydrogen envelope, and all the way to their end-of-life as cooling compact remnant white dwarfs (see the reviews by Chaplin & Miglio 2013; Hekker & Christensen-Dalsgaard 2017; Córsico et al. 2019; García & Ballot 2019; Giammichele et al. 2022). Multiple reasons for this focus on stars born with a low mass (defined here as stars with a mass in the range 0.7 M M⊙) are relevant. Low-mass stars have the following characteristics:
@@
They constitute the dominant population in our Milky Way galaxy, allowing the study of its history, archaeology, and current structure. This is currently being done from the multitude of Gaia data (Gaia Collaboration et al. 2016, 2023b) and ongoing or near-future dedicated all-sky survey high-resolution spectroscopy (Kollmeier et al. 2017; Pinsonneault et al. 2018; Bensby et al. 2019; Jin et al. 2023);
@@
They reveal high-frequency solar-like oscillations excited stochastically by the turbulent convection in their envelopes throughout almost their entire nuclear burning life. Such oscillations obey scaling relations, which were already summarised by Kjeldsen & Bedding (1995) long before the era of space asteroseismology. These scaling relations make it easy to deduce the mass, radius, and age from radial, dipole, and quadrupole modes, with high precisions of a few percent. This is achieved by comparing the observed properties of detected low-degree global oscillation modes with those of the Sun (see e.g. García & Ballot 2019, for an extensive review). It has the major advantage that one can use the heritage from helioseismology (Christensen-Dalsgaard 2002, 2021) and transfer it to asteroseismic applications to solar-like oscillations;
@@
They cover the most important parameter space of exoplanet host stars, including those with multiple Earth-like rocky planets in the habitable zone;
@@
They have a relatively modest binary and multiplicity fraction compared to stars of higher mass (see Fig. 2);
@@
They are all slow rotators from early in their life as they are subject to efficient magnetic braking during their core hydrogen burning phase. As we detail in the next sections, this implies that the impact of rotation on the modelling of their oscillations is in the easiest regime of the frequency domain. We come back to this important aspect throughout this review.
@@
None of these five aspects is valid for moderate to fast rotators of intermediate to high mass. We define stars born with 1.3 M M⊙ as intermediate-mass stars and those with M⊙ as high-mass stars. For the pulsators in these two mass regimes we do not have a main-sequence calibrator like the Sun, whose helioseismology (Christensen-Dalsgaard 2002) has been essential to guide similar applications to low-mass stars.
@@
Though a minority compared to their low-mass colleagues, stars born with M⊙ deliver a major fraction of the energy and heavy elements to galaxies and to the Universe as a whole. For this reason, and given the fast progress of asteroseismology for this under-represented category of fast rotators, we dedicate most of this review to asteroseismic applications of stars born with intermediate or high mass. In the next section we clarify what we mean by moderate and fast rotation.
@@
Figure 2 shows the binary fraction for stars and brown dwarfs. This figure makes it clear that we cannot ignore stellar multiplicity in a review focusing on the asteroseismology of intermediate- and high-mass stars. In particular, close binarity dominates the evolution of the most massive stars (Sana et al. 2012, 2013, 2014). While this was firmly established observationally a decade ago, it is recognised that binary evolution theory is still affected significantly by the major uncertainties occurring in the theory of single-star evolution (Marchant & Bodensteiner 2024). Hence, this must be given the highest priority, and asteroseismology is the optimal way to do so (Aerts 2021). We touch upon various aspects of binarity and multiplicity for asteroseismic studies in Sect. 5 and Sect. 6. The class of sub-dwarfs is discussed briefly in Sect. 7. It deserves special attention in this respect as the binary fraction is high (Han et al. 2003; Hu et al. 2008; Vos et al. 2017, 2019). In fact, all sub-dwarfs originate from binary interactions because the seemingly single class members can only be understood in terms of a common envelope phase resulting in a merger (Heber 2009). Due to the efficient loss of their envelope, the core-helium burning sub-dwarfs are of low mass in their current phase of evolution (about 0.5 M⊙). Those with oscillations reveal high-frequency modes with periods much shorter than their rotation and/or orbital period (see Lynas-Gray 2021, for a review). We also briefly touch upon merger seismology for higher mass objects, offering a way to hunt for such stars asteroseismically. Section 8 is dedicated to future outlooks, putting emphasis on Gaia’s role in asteroseismology.
@@
Rather than trying to be exhaustive in reviewing the progress of asteroseismology, this paper takes the approach of offering the non-expert reader highlights from the recent literature based on prominent case studies. This way, we illustrate what can be achieved with current and future data and theoretical developments. The plan of the paper is as follows. We provide an accessible introduction to asteroseismology in Sect. 3, touching upon the wave equations representing stellar oscillations and the accompanying frequency regimes. Section 4 covers recent applications of asteroseismology of single stars, starting with the Sun and gradually increasing the level of complexity in the mode computations as the ratio of the oscillation mode periods to the rotation period increases. In Sect. 5 we highlight how model-independent dynamical masses and radii help asteroseismology, before moving on to tidal asteroseismology in Sect. 6 and sub-dwarf and merger seismology in Sect. 7. We end the paper with an extensive outlook for the bright future of asteroseismic applications to stars of intermediate and high mass from combined Gaia data, survey spectroscopy, and ongoing TESS or future PLATO space photometry. There is a pertinent need for new theory and better modelling tools in order to interpret the oscillations of fast rotators and close binaries up to current measurement precisions.